3. ABGs at this time are: PaO2 110 mm Hg on 40% oxygen, PCO2 = 55 mm Hg, Bicarb = 26 mEq/L,

Calculate the pH!

4. What is your interpretation of the acid–base balance?

5. Which of the two scenarios is consistent with the primary problem?  

a. The subject had an LV MI, leading to decreased LV contractility and low C.O.

b. The subject had COPD and increased RV afterload leading to an RV MI and low C.O.
Answer #1

3)Henderson Hasselbach equayion

pH=6.1+log{10(HCO3÷0.0307 ×paCo2)}

=6.1+log {26÷0.0307 ×55}

=6.1+log { 46579.80)

=10.76

4) The result interprets that he has primary Metabolic alkalosis with superimposed Respiratory acidosis

5)This will cause increased RV after load and leading to RV MI and low CO